OpenStudy (anonymous):

Estimate the sum or difference. Use 0, 1/2, ans 1. 2/15+ 13/30 A 0 B 1/2 C 1

OpenStudy (elonasushchik):

B.

OpenStudy (michele_laino):

\[\frac{ 2 }{ 15 }+\frac{ 13 }{ 30 }=\frac{ 4 }{ 30 }+\frac{ 13 }{ 30 }=\frac{ 17 }{ 30 }\]